### v1

# ms-Mamba: Multi-scale Mamba for Time-Series Forecasting

Abstract
The problem of Time-series Forecasting is generally addressed by recurrent, Transformer-based and the recently proposed Mamba-based architectures. However, existing architectures generally process their input at a single temporal scale, which may be sub-optimal for many tasks where information changes over multiple time scales. In this paper, we introduce a novel architecture called Multi-scale Mamba (ms-Mamba) to address this gap. ms-Mamba incorporates multiple temporal scales by using multiple Mamba blocks with different sampling rates ($\Delta$s). Our experiments on many benchmarks demonstrate that ms-Mamba outperforms state-of-the-art approaches, including the recently proposed Transformer-based and Mamba-based models. Codes and models will be made available.
[…]
XXX
1 Introduction
Time-series Forecasting (TSF) is the problem of predicting future values of a variable of interest, given its history. This fundamental problem used to be generally addressed using recurrent architectures [37, 11] and long-short term memory [19] or their variants [7, 14], see, e.g., [24, 27] for detailed surveys. Such models are inherently well-suited to the task due to their sequential information modeling abilities. The introduction of self-attention based architectures, a.k.a. Transformers [32], enabled attending to more informative patterns and correlations across time and provided significant improvements. However, Transformers’ quadratic computational complexity has been a limiting factor.
State Space Models (SSMs) [17, 31] are reported to provide a better balance between performance and computational complexity. Recently proposed architectures based on SSMs, namely, Mamba [16], offer the promise of on-par or better performance than Transformer-based alternatives while running significantly faster. This has led to the widespread use of Mamba or its derivatives across different domains [16, 43, 30, 41].
The use of a Mamba-based approach for TSF was recently explored by Wang et al.[34]. Wang et al.introduced an architecture, called S-Mamba, which used Mamba in both the forward and reverse directions for TSF. Wang et al.showed that this simple approach obtained state-of-the-art (SOTA) results on many TSF benchmarks, often providing significant gains over Transformer-based architectures.
Figure 1: (a) Mamba and its variations (S-Mamba) use a single time-scale while processing time-series data. (b) Our ms-Mamba processes its input at different time-scales to better capture signal at different scales.
Time-series data generally consist of signals of multiple temporal scales. To better capture and exploit the multi-scale nature of time-series data, the literature has introduced extensions over the conventional models; e.g., multi-scale recurrent architectures [8], multi-scale convolution [22] and multi-scale Transformers [47, 5].
In this paper, we introduce ms-Mamba, a Mamba-based architecture for multi-scale processing of time-series data. To be specific, by leveraging on the versatility of SSMs’ learnable sampling rate, we construct a block that consists of multiple SSMs with different independent or inter-related sampling rates. We show that our ms-Mamba performs better than Transformer-based and Mamba-based architectures on several datasets.
Contributions. Our main contributions are as follows:
  • We propose a multi-scale architecture based on Mamba. To do so, we use multiple SSMs with different sampling rates to process the signal at different temporal scales.
  • We introduce and compare different strategies for using different sampling rates for different SSMs: (1) Using hyper-parameters as multipliers for a learned sampling rate, (2) learning different sampling rate for each SSM, (3) and estimating sampling rates from the input.
  • We show that, on the commonly used TSF benchmarks, our ms-Mamba surpasses or performs on par with SOTA models. For example, on the Solar-Energy dataset, ms-Mamba outperforms its closest competitor s-Mamba (0.229 vs. 0.240 in terms of mean-squared error) with less parameters, less memory footprint and less computational overhead.
2 Related Work
2.1 Time-series Forecasting
2.1.1 Transformer-based Models
Transformers, initially introduced by Vaswani et al.[32], have revolutionized tasks that involve sequence processing and generation, with their self-attention mechanism proving highly effective in capturing long-range dependencies. This architecture, originally designed for natural language processing, has since been adapted to time-series forecasting tasks [1], primarily due to its ability to model complex temporal relationships. Duong-Trung et al.[10] demonstrate the efficacy of Transformers in long-term multi-horizon forecasting, addressing the challenge of vanishing correlations over extended horizons.
Recent studies have aimed to address the limitations of standard Transformers in time-series applications. Foumani et al.[12] propose enhanced positional encodings to improve the positional awareness of the Transformer architecture in multivariate time-series classification. Lim et al.[25] propose a Transformer architecture to make use of a complex mix of inputs. Wang et al.[33] present Graphformer, a model that replaces traditional convolutional layers with dilated convolutional layers, thereby improving the efficiency of capturing temporal dependencies across multiple variates in a graph-based framework.
Despite their unprecedented successes in natural language processing tasks, Transformer models face several challenges when applied to other time-series domains. One key limitation is their content-based attention mechanism, which struggles to detect crucial temporal dependencies, particularly in cases where dependencies weaken over time or when strong seasonal patterns are present [38]. Additionally, Transformers suffer from the quadratic complexity of the attention mechanism, which increases computational costs and memory usage, for long input sequences [36].
To address these issues, several studies have proposed modifications to the self-attention mechanism. For instance, Zhou et al.[48] introduce Informer that employs a sparsified self-attention operation to lower computational complexity and improve long-term forecasting efficiency. Similarly, Wu et al.[39] propose Autoformer, which relies on an auto-correlation-based self-attention to better capture temporal dependencies.
2.1.2 Linear Models
Linear models are another popular approach in TSF due to their simplicity and efficiency [2]. Oreshkin et al.[29] propose a stacked MLP based architecture with residual links. Challu et al.[3] improve this architecture with multi-rate data sampling and hierarchical interpolation for effectively modeling extra long sequences. Zeng et al.[44] analyze Transformers for TSF and found that simple linear mappings can outperform Transformer models especially when the data has strong periodic patterns. Chen et al.[6] introduce another notable linear approach, TSMixer, which leverages an all-MLP architecture to efficiently incorporate cross-variate and auxiliary information. Zhang et al.[45] propose LightTS which is tailored towards efficiently handling very long input series in multivariate TSF. Wang et al.[35] propose time-series Multi-layer Perceptron (MLP), which improves forecasting performance by incorporating domain-specific knowledge into the MLP architecture.
While linear models with MLPs are simpler architectures and faster compared to Transformer-based models, they face several limitations. These models generally struggle with non-linear dependencies and tend to underperform in scenarios involving highly volatile or non-stationary patterns [6]. Moreover, compared to Transformer-based models, linear architectures are less effective at capturing global dependencies. This limitation necessitates longer input sequences to achieve comparable forecasting performance, which can increase the computational cost [42].
2.2 Mamba Models
Mamba [16] is a recent sequence model based on State Space Models (SSMs) [17, 31]. Due to its promise of better efficiency-performance trade-off, Mamba has quickly attracted interest from researchers across different domains [30, 41]. Mamba’s ability to perform content-based reasoning in linear complexity to the sequence length with its hardware-aware algorithm, made it an attractive alternative to the Transformer models. Several works have explored its application to time-series forecasting tasks. Wang et al.[34] propose S-Mamba, which relies on a bidirectional Mamba layer to capture inter-variate dependencies and an MLP to extract temporal dependencies. Their model achieves SOTA performance while being faster than Transformer-based alternatives.
Figure 2: An overview of the proposed method. ms-Mamba processes the time-series data at different sampling rates to better capture the multi-scale nature of the input signal. This is achieved by processing and updating the embeddings with different sampling rates (SR).
3 Preliminaries and Background
3.1 Problem Formulation
Time-series forecasting is the problem of estimating future $F$ values $\mathbf{Y}_{t+1:t+F}=\{\mathbf{x}_{t+1},\mathbf{x}_{t+2},...,\mathbf{x}_{t+F}\}\in\mathbb{R}^{F\times D}$ of a multi-variate time-series data given its recent $L$ values $\mathbf{X}_{t-L:t}=\{\mathbf{x}_{t-L},\mathbf{x}_{t-L+1},...,\mathbf{x}_{t}\}\in\mathbb{R}^{L\times D}$ as input. The task is to find the mapping $f$:
$f:\mathbf{X}_{t-L:t}\rightarrow\mathbf{Y}_{t+1:t+F},$ (1)
which is represented by a deep network whose parameters are estimated from a training dataset.
3.2 State Space Models (SSMs)
SSMs [17, 31] are sequence models which use a latent state space representation for representing a mapping between a time-series input $x(t)$ and the output $y(t)$ (considering a single-variate setting to simplify notation):
$\delta h(t)$ $=$ $\textbf{A}h(t)+\textbf{B}x(t),$ (2)
$y(t)$ $=$ $\textbf{C}h(t),$ (3)
where $\mathbf{A},\mathbf{B},\mathbf{C}$ are matrices with learnable values; $h(t)$ is the latent (state) representation; and $\delta h(t)$ is the update for the latent space with the current input, $x(t)$. This continuous formulation is transformed into a discrete model with a sampling rate $\Delta$ as follows:
$h_{t}$ $=$ $\hat{\textbf{A}}h_{t-1}+\hat{\textbf{B}}x_{t},$ (4)
$y_{t}$ $=$ ${\textbf{C}}h_{t},$ (5)
where $\hat{\textbf{A}},\hat{\textbf{B}},{\textbf{C}}$ are derived by the chosen sampling function (e.g., for zeroth-order hold sampling, $\hat{\textbf{A}}=\exp(\Delta\mathbf{A})$, $\hat{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A}-I)\cdot\Delta\mathbf{B}$ [16]).
SSMs have been recently extended to work more efficiently through the use of a convolutional approach [18] and more effectively through better initialization [13]. Moreover, by relating $\hat{\textbf{A}},\hat{\textbf{B}},{\textbf{C}}$ to the input, the Mamba model [16] provides comparable or better performance than its Transformer-based counterparts.
4 Methodology: ms-Mamba
In this section, we describe our ms-Mamba in detail. For multi-scale temporal processing, ms-Mamba essentially leverages multiple Mamba blocks with different $\Delta$ working in parallel. See Figure 2 for an overview.
4.1 Embedding Layer
Following prior work [15, 26], we first transform the input time-series data through an embedding layer (Figure 2). Given an input sequence $\mathbf{X}\in\mathbb{R}^{L\times D}$ with $L$ time steps and $D$ variables, we apply a linear transformation along the temporal dimension:
$\mathbf{E}=\mathrm{Embedding}(\mathbf{X})\in\mathbb{R}^{D_{e}\times D},$ (6)
where $D_{e}$ is the embedding dimension. This transformation maps each time-series from length $L$ to length $D_{e}$ while preserving the number of variables $D$, thus enabling us to deal with fixed-length tokens instead of variable input sequence length $L$.
4.2 Multi-scale Mamba Layer
As summarized in Section 3.2, SSMs, Mamba and their variants process time at one learnable sampling rate, $\Delta$. Our architecture ms-Mamba addresses this gap by processing the input at different sampling rates $\Delta_{1}$, $\Delta_{2}$, …, $\Delta_{s}$. This is achieved by combining multiple Mamba blocks with different sampling rates as follows:
$\mathbf{E}^{l}_{m}=\mathrm{Avg}(\mathrm{Mamba}(\mathbf{E}^{l};\Delta_{1}),...,\mathrm{Mamba}(\mathbf{E}^{l};\Delta_{n})),$ (7)
where $\mathbf{E}^{l}$ is the output of the embedding layer at layer $l$. We explore three different strategies for obtaining $\Delta_{i}$:
  1. 1.
    Fixed temporal scales, where $\Delta_{1}$ is kept learnable (as in the original Mamba model) but $\Delta_{2},\Delta_{3},...,\Delta_{n}$ are taken as multiples of $\Delta_{1}$:
    $\Delta_{i}=\alpha_{i}\times\Delta_{1},\quad\quad i\in\{2,...,n\},$ (8)
    where $\alpha_{i}$ are hyper-parameters.
  2. 2.
    Learnable temporal scales, where all $\Delta_{i}$ are defined as learnable variables as in the original Mamba model.
  3. 3.
    Dynamic temporal scales, where all $\Delta_{i}$ are estimated through a Multi-layer Perceptron:
    $\Delta_{i}=\mathrm{MLP}(\mathrm{Flatten}(\mathbf{E}^{l})),$ (9)
    where $\mathrm{Flatten}(\cdot)$ reshapes the input tensor $\mathbf{E}^{l}\in\mathbb{R}^{L\times D_{e}}$ into a vector of dimension $L\cdot D_{e}$, and $\mathrm{MLP}(\cdot)$ consists of two linear layers with a ReLU activation in between: $\mathrm{MLP}(\mathbf{x})=\mathbf{W}_{2}\max(0,\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}$, mapping the flattened input to $n$ different sampling rates.
To improve the effectiveness of sequential processing, we employ our Multi-scale Mamba module in both directions as illustrated in Figure 2, following prior work (e.g., [15, 50]).
4.3 Normalization, Feed-Forward Network and Projection
The output of the Multi-scale Mamba Layer ($\mathbf{E}^{l}_{m}$) passes through Layer Normalization, a multi-layer perceptron (MLP – with two layers with the ReLU nonlinearity) to obtain the embeddings for the next layer ($\mathbf{E}^{l+1}\in\mathbb{R}^{L\times D_{e}}$):
$\mathbf{E}^{l+1}=\mathrm{MLP}(\mathrm{LayerNorm}(\mathbf{E}^{l}_{m})).$ (10)
After the last encoder block ($\mathbf{E}^{L}\in\mathbb{R}^{L\times D_{e}}$), a linear projection layer is applied to map the embedding dimension to prediction length to obtain the final prediction ($\hat{y}\in\mathbb{R}^{F\times D}$):
$\hat{y}=\mathrm{Linear(\mathbf{E}^{L})}.$ (11)
4.4 Training Objective
The model is trained to minimize the Mean Square Error (MSE) between the predicted values and the ground truth:
$\mathcal{L}=\frac{1}{F\times D}\sum_{i=1}^{F}\sum_{j=1}^{D}(\hat{y}_{i,j}-y_{i,j})^{2},$ (12)
where $\hat{y}_{i,j}$ and $y_{i,j}$ are the predicted and ground truth values for the $i$-th time step and $j$-th variable, respectively; $F$ is the forecast horizon; and $D$ is the number of variables. The model parameters are optimized using the Adam optimizer [20] – see the Suppl. Mat. for more details about the training and experimental details.
5 Experiments
5.1 Experimental Details
Datasets. To evaluate our proposed ms-Mamba, we conduct extensive experiments on thirteen real-world time-series forecasting benchmark datasets. The datasets are grouped into three categories for easier comparison. (1) Traffic-related datasets, which include Traffic [39] and PEMS [4]. The Traffic dataset consists of hourly road occupancy rates from the California Department of Transportation, consisting of data collected from 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016. PEMS datasets are complex spatial-temporal datasets for California’s public traffic networks, includes four subsets (PEMS03, PEMS04, PEMS07, PEMS08), similar to SCINet. These traffic-related datasets has many periodic features. (2) ETT (Electricity Transformer Temperature) datasets [48], which contain load and oil temperature data from electricity transformers, collected between July 2016 and July 2018. This group includes four subsets: ETTm1, ETTm2, ETTh1, and ETTh2, which have fewer variables and show less regularity compared to traffic datasets. (3) Other datasets: Electricity [39], Exchange [39], Weather [39], and Solar-Energy [21]. The Electricity dataset includes the hourly electricity usage of 321 customers from 2012 to 2014. Solar-Energy dataset contains solar power generation data from 137 PV plants in Alabama in 2006, recorded at 10 minute resolution. The Weather dataset includes 21 meteorological indicators also recorded at 10 minute resolution from the Max Planck Biogeochemistry Institute’s Weather Station in 2020. Exchange dataset compiles daily exchange rates for eight countries from 1990 to 2016. The prior two datasets of this category contain many features most of which are periodic, the last two datasets have fewer primarily aperiodic features. See the Suppl. Mat for a summary of the datasets.
Models ms-Mamba ms-Mamba w/ Learnable $\Delta$ S-Mamba iTrans- former RLinear PatchTST Cross- former TiDE TimesNet DLinear FED- former Auto- former
Traffic 96 0.376 0.375 0.382 0.395 0.649 0.462 0.522 0.805 0.593 0.650 0.587 0.613
192 0.392 0.384 0.396 0.417 0.601 0.466 0.530 0.756 0.617 0.598 0.604 0.616
336 0.405 0.408 0.417 0.433 0.609 0.482 0.558 0.762 0.629 0.605 0.621 0.622
720 0.452 0.442 0.460 0.467 0.647 0.514 0.589 0.719 0.640 0.645 0.626 0.660
Avg 0.406 0.402 0.414 0.428 0.626 0.481 0.550 0.760 0.620 0.625 0.610 0.628
PEMS03 12 0.065 0.066 0.065 0.071 0.126 0.099 0.090 0.178 0.085 0.122 0.126 0.272
24 0.087 0.087 0.087 0.093 0.246 0.142 0.121 0.257 0.118 0.201 0.149 0.334
48 0.131 0.133 0.133 0.125 0.551 0.211 0.202 0.379 0.155 0.333 0.227 1.032
96 0.197 0.201 0.201 0.164 1.057 0.269 0.262 0.490 0.228 0.457 0.348 1.031
Avg 0.120 0.122 0.122 0.113 0.495 0.180 0.169 0.326 0.147 0.278 0.213 0.667
PEMS04 12 0.074 0.072 0.076 0.078 0.138 0.105 0.098 0.219 0.087 0.148 0.138 0.424
24 0.086 0.083 0.084 0.095 0.258 0.153 0.131 0.292 0.10 0.224 0.177 0.459
48 0.102 0.099 0.115 0.120 0.572 0.229 0.205 0.409 0.136 0.355 0.270 0.646
96 0.130 0.121 0.137 0.150 1.137 0.291 0.402 0.492 0.190 0.452 0.341 0.912
Avg 0.098 0.094 0.103 0.111 0.526 0.195 0.209 0.353 0.129 0.295 0.231 0.610
PEMS07 12 0.060 0.060 0.063 0.067 0.118 0.095 0.094 0.173 0.082 0.115 0.109 0.199
24 0.075 0.075 0.081 0.088 0.242 0.150 0.139 0.271 0.101 0.210 0.125 0.323
48 0.091 0.091 0.093 0.110 0.562 0.253 0.311 0.446 0.134 0.398 0.165 0.390
96 0.111 0.109 0.117 0.139 1.096 0.346 0.396 0.628 0.181 0.594 0.262 0.554
Avg 0.085 0.084 0.089 0.101 0.504 0.211 0.235 0.380 0.124 0.329 0.165 0.367
PEMS08 12 0.074 0.073 0.076 0.079 0.133 0.168 0.165 0.227 0.112 0.154 0.173 0.436
24 0.097 0.098 0.104 0.115 0.249 0.224 0.215 0.318 0.141 0.248 0.210 0.467
48 0.156 0.154 0.167 0.186 0.569 0.321 0.315 0.497 0.198 0.440 0.320 0.966
96 0.243 0.236 0.245 0.221 1.166 0.408 0.377 0.721 0.320 0.674 0.442 1.385
Avg 0.143 0.140 0.148 0.150 0.529 0.280 0.268 0.441 0.193 0.379 0.286 0.814
Table 1: Experiment 1: Quantitative comparison between ms-Mamba and the existing methods on traffic-related datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 12, 24, 48, 96 for PEMS and 96, 192, 336, 720 for Traffic. Top results are highlighted in bold while the second bests are underlined.
Compared Methods. We compare our model with 10 state-of-the-art (SOTA) time-series forecasting models belonging to 4 different model families: (1) Mamba based models: S-Mamba [34]; (2) Transformer based models: iTransformer [26], PatchTST [28], Crossformer [46], FEDformer [49], Autoformer [39]; (3) Linear based models: TiDE [9], DLinear [44], RLinear [23]; and (4) Temporal Convolution based models: TimesNet [40]. The following provides a brief overview of these models:
  • S-Mamba [34] employs a bidirectional Mamba encoder block to capture inter-variate correlations and a feed forward network temporal dependency encoding layer to learn temporal sequence dynamics. This novel approach is the current SOTA model for TSF task and forms the foundation of our proposed method.
  • iTransformer [26] inverts the order of sequence processing by first analyzing each individual variate separately and then merging the information across all variates.
  • PatchTST [28] divides the time-series into sub-series patches treated as input tokens. It leverages channel-independent shared embeddings and weights for efficient representation learning.
  • Crossformer [46] embeds multivariate time-series into a 2D vector array preserving time and dimension information and introduces two-stage attention to capture both cross-time and cross-dimension dependencies.
  • FEDformer [49] is a frequency-enhanced Transformer that uses trend-seasonality decomposition and exploit sparse representations, Fourier transform, of time-series data to achieve linear complexity to sequence length.
  • Autoformer [39] constructs a decomposition architecture that employs traditional sequence decomposition in its inner blocks and utilizes an auto-correlation mechanism.
  • DLinear [44] maps trend and seasonality components into predictions via a single linear layer.
  • TiDE [9] is a MLP based encoder-decoder model that is best suitable for linear dynamical systems.
  • RLinear [23] is the current SOTA linear model that introduces reversible normalization and channel independence within a purely linear structure.
  • TimesNet [40] proposes a task-general backbone, TimesBlock, that transforms 1D time-series into 2D tensors and uses 2D convolution kernels to capture intra-period and inter-period variations.
Training and Implementation Details. See the Suppl Mat for training and implementation details, especially the hyperparameters.
Performance Measure. Following the common practice (e.g., [34, 26, 28]), models’ performances are compared using the Mean Square Error (MSE) as defined in Equation (12).
Models ms-Mamba ms-Mamba w/ Learnable $\Delta$ S-Mamba iTrans- former RLinear PatchTST Cross- former TiDE TimesNet DLinear FED- former Auto- former
ETTm1 96 0.328 0.326 0.333 0.334 0.355 0.329 0.404 0.364 0.338 0.345 0.379 0.505
192 0.372 0.372 0.376 0.377 0.391 0.367 0.450 0.398 0.374 0.380 0.426 0.553
336 0.406 0.406 0.408 0.426 0.424 0.399 0.532 0.428 0.410 0.413 0.445 0.621
720 0.470 0.470 0.475 0.491 0.487 0.454 0.666 0.487 0.478 0.474 0.543 0.671
Avg 0.394 0.394 0.398 0.407 0.414 0.387 0.513 0.419 0.400 0.403 0.448 0.588
ETTm2 96 0.176 0.175 0.179 0.180 0.182 0.175 0.287 0.207 0.187 0.193 0.203 0.255
192 0.244 0.244 0.250 0.250 0.246 0.241 0.414 0.290 0.249 0.284 0.269 0.281
336 0.309 0.306 0.312 0.311 0.307 0.305 0.597 0.377 0.321 0.369 0.325 0.339
720 0.408 0.407 0.411 0.412 0.407 0.402 1.730 0.558 0.408 0.554 0.421 0.433
Avg 0.284 0.283 0.288 0.288 0.286 0.281 0.757 0.358 0.291 0.350 0.305 0.327
ETTh1 96 0.384 0.384 0.386 0.386 0.386 0.414 0.423 0.479 0.384 0.386 0.376 0.449
192 0.437 0.438 0.443 0.441 0.437 0.460 0.471 0.525 0.435 0.436 0.420 0.500
336 0.479 0.482 0.489 0.487 0.479 0.501 0.570 0.565 0.491 0.481 0.459 0.521
720 0.482 0.493 0.502 0.503 0.481 0.500 0.653 0.594 0.521 0.519 0.506 0.514
Avg 0.445 0.449 0.455 0.454 0.446 0.469 0.529 0.541 0.458 0.456 0.440 0.496
ETTh2 96 0.291 0.291 0.296 0.297 0.288 0.302 0.745 0.400 0.340 0.333 0.358 0.346
192 0.371 0.369 0.376 0.380 0.374 0.388 0.877 0.528 0.402 0.477 0.429 0.456
336 0.411 0.412 0.424 0.428 0.415 0.426 1.043 0.643 0.452 0.594 0.496 0.482
720 0.418 0.418 0.426 0.427 0.420 0.431 1.104 0.874 0.462 0.831 0.463 0.515
Avg 0.373 0.373 0.381 0.383 0.374 0.387 0.942 0.611 0.414 0.559 0.437 0.450
Table 2: Experiment 1: Quantitative comparison between ms-Mamba and the existing methods on ETT Datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720. Top results are highlighted in bold while the second bests are underlined.
Models ms-Mamba ms-Mamba w/ Learnable $\Delta$ S-Mamba iTrans- former RLinear PatchTST Cross- former TiDE TimesNet DLinear FED- former Auto- former
Electricity 96 0.137 0.138 0.139 0.148 0.201 0.181 0.219 0.237 0.168 0.197 0.193 0.201
192 0.157 0.157 0.159 0.162 0.201 0.188 0.231 0.236 0.184 0.196 0.201 0.222
336 0.171 0.174 0.176 0.178 0.215 0.204 0.246 0.249 0.198 0.209 0.214 0.231
720 0.195 0.199 0.204 0.225 0.257 0.246 0.280 0.284 0.220 0.245 0.246 0.254
Avg 0.165 0.167 0.170 0.178 0.219 0.205 0.244 0.251 0.192 0.212 0.214 0.227
Exchange 96 0.085 0.086 0.086 0.086 0.093 0.088 0.256 0.094 0.107 0.088 0.148 0.197
192 0.177 0.178 0.182 0.177 0.184 0.176 0.470 0.184 0.226 0.176 0.271 0.300
336 0.325 0.326 0.332 0.331 0.351 0.301 1.268 0.349 0.367 0.313 0.460 0.509
720 0.843 0.843 0.867 0.847 0.886 0.901 1.767 0.852 0.964 0.839 1.195 1.447
Avg 0.358 0.358 0.367 0.360 0.378 0.367 0.940 0.370 0.416 0.354 0.519 0.613
Weather 96 0.163 0.163 0.165 0.174 0.192 0.177 0.158 0.202 0.172 0.196 0.217 0.266
192 0.213 0.213 0.214 0.221 0.240 0.225 0.206 0.242 0.219 0.237 0.276 0.307
336 0.269 0.270 0.274 0.278 0.292 0.278 0.272 0.287 0.280 0.283 0.339 0.359
720 0.349 0.349 0.350 0.358 0.364 0.354 0.398 0.351 0.365 0.345 0.403 0.419
Avg 0.249 0.249 0.251 0.258 0.272 0.259 0.259 0.271 0.259 0.265 0.309 0.338
Solar Ener. 96 0.196 0.195 0.205 0.203 0.322 0.234 0.310 0.312 0.250 0.290 0.242 0.884
192 0.230 0.230 0.237 0.233 0.359 0.267 0.734 0.339 0.296 0.320 0.285 0.834
336 0.250 0.247 0.258 0.248 0.397 0.290 0.750 0.368 0.319 0.353 0.282 0.941
720 0.249 0.249 0.260 0.249 0.397 0.289 0.769 0.370 0.338 0.356 0.357 0.882
Avg 0.231 0.229 0.240 0.233 0.369 0.270 0.641 0.347 0.301 0.330 0.291 0.885
Table 3: Experiment 1: Quantitative comparison between ms-Mamba and the existing methods on Electricity, Exchange, Weather and Solar-Energy Datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720. Top results are highlighted in bold while the second bests are underlined.
5.2 Experiment 1: Comparison with State-of-the-Art
We compare ms-Mamba with fixed and learnable temporal scales against SOTA methods over 13 benchmark datasets.
Traffic-related Datasets (Table 1). The results in Table 1 show that ms-Mamba with fixed or learnable temporal scales provides the best or second best results over all traffic datasets across all forecast lengths. Compared to our baseline model of S-Mamba, ms-Mamba delivers significant improvements, especially on the Traffic dataset.
ETT Datasets (Table 2). On ETT datasets, as in Table 2, ms-Mamba with fixed or learnable temporal scales is typically the second best and the best performing method in ETTh2 and in some configurations of ETTm1 and ETTm2. Compared to S-Mamba, ms-Mamba provides better performance.
Other Datasets (Table 3). The results in other datasets (Table 3) are in agreement with the results on Traffic-related and ETT datasets: our ms-Mamba with fixed or learnable temporal scales performs better than or is generally on par with SOTA methods, and consistently provides better performance than the S-Mamba baseline.
                                    Dataset $\Rightarrow$ Traffic ETTh1 Solar Energy
                             VC Encoding $\Rightarrow$ 96 192 336 720 96 192 336 720 96 192 336 720
ms-Mamba with fixed scales: $\mathbf{\alpha}=(1,2,4,8)$ 0.376 0.392 0.405 0.452 0.384 0.437 0.479 0.482 0.196 0.230 0.250 0.249
ms-Mamba with fixed scales: $\mathbf{\alpha}=(0.5,1,1.5,2)$ 0.374 0.389 0.414 0.458 0.386 0.438 0.481 0.491 0.197 0.232 0.248 0.251
ms-Mamba with fixed scales: $\mathbf{\alpha}=(1,2,3,4)$ 0.390 0.403 0.415 0.455 0.386 0.439 0.482 0.495 0.196 0.232 0.250 0.251
ms-Mamba with fixed scales: $\mathbf{\alpha}=(1,4,8,16)$ 0.380 0.411 0.421 0.453 0.385 0.439 0.484 0.495 0.197 0.232 0.250 0.251
ms-Mamba with learnable scales 0.375 0.384 0.408 0.442 0.384 0.438 0.482 0.493 0.195 0.230 0.247 0.249
ms-Mamba with dynamic scales 0.376 0.390 0.414 0.454 0.384 0.440 0.480 0.493 0.194 0.231 0.249 0.251
Table 4: Experiment 2: Ablation study on Traffic, ETTh1 and Solar Energy datasets (one dataset from each category). The lookback length $L=96$, while the forecast length $T\in\left\{96,192,336,720\right\}$. $\mathbf{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})$ indicates that the $\Delta_{1}$ (learnable sampling rate of the base Mamba) is multiplied by these coefficients to obtain the sampling rates for the Mamba blocks. Top results are highlighted in bold while the second bests are underlined.
5.3 Experiment 2: Ablation Analysis
In this experiment, we investigate the different strategies for integrating multiple-scales into ms-Mamba. For this, we evaluate the performances of the strategies described in Section 4.2: (i) ms-Mamba with fixed temporal scales, where we multiply the learnable sampling rate $\Delta_{1}$ by fixed hyperparameters, $\Delta_{i}=\alpha_{i}\times\Delta_{1}$. (ii) ms-Mamba with learnable temporal scales, where each $\Delta_{i}$ is a learnable parameter. (iii) ms-Mamba with dynamic temporal scales, where each $\Delta_{i}$ is estimated by an MLP applied on the input embeddings. For this analysis, we tune the hyperparameters in all settings. To keep the number of experiments manageable, we consider one dataset from each category.
The results in Table 4 suggest that using fixed temporal scales with $\alpha_{i}$ coefficients of $(1,2,4,8)$ performs best among different coefficients considered (6 best and 2 second-best results out of 12 datasets). Smaller or larger coefficients than $(1,2,4,8)$ do not appear to provide better results overall. Moreover, ms-Mamba with learnable $\Delta_{i}$ provides slightly better results than the best fixed temporal scale version (provides 6 best and 4 second-best results).
Fixed temporal scales introduces several hyperparameters to be tuned, which is a significant limitation. Therefore, ms-Mamba with learnable temporal scales can be the preferred option. However, in Experiment 1, we have considered both versions for a more extensive evaluation.
Models ms-Mamba S-Mamba
Metric MSE #Params Memory MACs MSE #Params Memory MACs
ETTh2 96 0.291 0.481M 1.84MB 0.165G 0.296 1.150M 4.40MB 1.563G
192 0.369 0.484M 1.85MB 0.171G 0.376 1.175M 4.48MB 1.580G
336 0.412 0.503M 1.92MB 0.180G 0.424 1.212M 4.62MB 1.606G
720 0.418 0.552M 2.11MB 0.195G 0.426 1.311M 5.00MB 1.675G
Solar Ener. 96 0.195 3.958M 15.10MB 16.72G 0.205 4.643M 17.71MB 20.00G
192 0.230 2.028M 7.74MB 8.57G 0.237 4.692M 17.90MB 20.21G
336 0.247 4.015M 15.31MB 16.99G 0.258 4.766M 18.18MB 20.54G
720 0.249 4.113M 15.70MB 17.43G 0.260 4.963M 18.93MB 21.40G
Traffic 96 0.375 29.68M 113.2MB 403.5G 0.382 9.186M 35.04MB 125.0G
192 0.384 14.94M 56.99MB 203.1G 0.396 9.236M 35.23MB 125.7G
336 0.408 29.81M 113.7MB 405.2G 0.417 9.310M 35.51MB 126.7G
720 0.442 29.68M 114.5MB 407.9G 0.460 9.507M 36.26MB 129.5G
Table 5: Experiment 3: Performance comparison of ms-Mamba and S-Mamba on one dataset from each category. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720.
5.4 Experiment 3: Efficiency Analysis
In this experiment, we investigate the efficiency of ms-Mamba (with learnable temporal scales, as it provides the best results and incurs more learnable parameters) in comparison with the baseline S-Mamba. In each dataset and for each method, we provide the results of the best performing configurations.
As listed in Table 5, on the ETTh1 and Solar Energy datasets, we see that ms-Mamba provides the best results with less parameters, memory and operations. This is crucial as it shows multiple temporal scales can be utilized with less computational overhead. However, this result is not observed in the Traffic dataset because it contains significantly more variates (862) compared to the ETTh2 (7) and Solar Energy (137) datasets. Although ms-Mamba provides better results with more variates, it is not able to do so with less parameters, memory and computations as in ETTh2 and Solar Energy datasets.
6 Conclusion
In this paper, we introduce a novel multi-scale architecture for the problem of time-series forecasting (TSF). Our architecture extends Mamba (or its derivative S-Mamba) where we include several Mamba blocks with different sampling rates to process multiple temporal scales simultaneously. The different sampling rates can either be fixed or learned from the data, which leads to a simple architecture with a multi-scale ability. Our results on several TSF benchmarks show that our approach provides the best or on par performance compared to SOTA methods. What is remarkable is that, compared to the baseline model (S-Mamba), ms-Mamba provides better results with less parameters, memory, and operations.
Limitations and Future Work. It is a promising research direction to apply ms-Mamba for other types of modalities, e.g., text and images. Moreover, ms-Mamba can complement other types of deep modules, e.g., scaled-dot-product attention, linear attention.